OpenStudy (anonymous):

Write the equation y - 2 = 4(x + 5) in standard form.

OpenStudy (anonymous):

y - 2 = 4x + 20 y = 4x + 22

OpenStudy (slaaibak):

standard form: y = mx + C Now multiply and simplify

OpenStudy (saifoo.khan):

\[\checkmark\]